Consider taking one sample from the distribution N(0,02). Is |x| a sufficient statistic?
Answer #1

The probability density function of one sample from N(0,02) is,

f(0:0) = 1200 e/20 11:0) =

5. 2πσ e-/272

= o(u(2):0) xh(2)

where

$(u(-2));0) = -11° /20

and

har) = 1

Thus, by factorization theorem and $(2:0) = o(u :0) x h(x) for all x and \sigma.

|x| is a sufficient statistic for \sigma.